Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 119     Entry time: 07 Jun 2024, 23:12
Author: Chan Kim 
Group: SPB 
Subject: Shift summary 

X-ray delivery
- 7.0 keV, up to 202 pulses at 564 kHz, 1.0 mJ.
- Average pulse energy was about 1.9 mJ on Tuesday and Wednesday. It has to be discussed with BKR tomorrow morning.

Optical laser delivery
- ns laser for jet illumination

Achievements/Observations
- BG measurement with water, buffer low salt
- Lysozyme measurement for geometry file
- Data collection for six different sample concentrations [2 x 15 minutes runs]

Issues
- Pulse energy is very low as compared with Tuesday and Wednesday in this week.
- Damage found on IRU_SLIT4 Top and Bottom blades.
- NKB focus moved by ~20um at around 15:00 (during the handover period). SideMic X moved +20um to compensate it.

ELOG V3.1.4-7c3fd00